OpenStudy (anonymous):

I'm stuck on these two questions in Precalc. Please help? I have the answers, but I would like to know how to get them. 1. (tan^3x-1)/(tanx-1)-sec^2x-1=0, cotx=? [answer is 1] 2. x=1/2tanx where -pi/2

OpenStudy (anonymous):

first problem in equation editor ..?? plz and what do have to find in ques.2

OpenStudy (anonymous):

\[(\tan^3\theta-1/\tan \theta -1)-\sec^2\theta -1=0\]

OpenStudy (anonymous):

For the second one, I'm supposed to write f(x) in terms of a single trigonometric function of theta.

OpenStudy (anonymous):

\[\frac{ (tanx-1)(\tan^2x+1+tanx )}{ tanx-1} -(1+\tan^2x)-1\]

OpenStudy (anonymous):

to make that fast i used x instead of theta ...so as u can see ..apply the identity of \[a^3-b^3\]

OpenStudy (anonymous):

so cotx =1 after simplifying

OpenStudy (anonymous):

That makes sense, thanks.

OpenStudy (anonymous):

in the second one ...does\[x=\frac{ 1 }{ 2tanx }\]

OpenStudy (anonymous):

\[x=1/2(tanx)\] is that the same thing?

OpenStudy (anonymous):

because by that my answer is coming as 1/2secx

OpenStudy (anonymous):

\[f(x)=\frac{ \frac{ 1 }{ 2tanx } }{ \sqrt{1+\frac{ 4 }{ 4\tan^2x }} }\]

OpenStudy (anonymous):

\[f(x)=\frac{ \frac{ 1 }{ 2tanx} }{ \frac{ secx }{ tanx } }\]

OpenStudy (anonymous):

hope u r getting that .??

OpenStudy (anonymous):

Yes, I think so.

OpenStudy (anonymous):

so hence ..the answer is 1/2secx ..??

OpenStudy (anonymous):

Maybe the top of the equation should be 2/tanx? x is supposed to be (1/2)(tanx). Does that make it come out to 1/2sinx?

OpenStudy (anonymous):

yes then it comes that .!

OpenStudy (anonymous):

Okay, I think I've got it now :) Thanks for your help!